The number of bacteria in a culture is given by b(t) = 1400e^0.25t, where t is measured in hours. After how many hours will the number of bacteria in the culture be three times that of the initial number of bacteria?
Answer by Cromlix(4381)

b(t) = 1400e^0.25t,
At t = 0 the number of bacteria
will be = 1400
3 * 1400 = 4200
b(t) = 1400e^0.25t,
4200 = 1400 e^0.25t
4200/1400 = e^0.25t
3 = e^0.25t
loge 3 = 0.25t loge e (loge e = 1)
t = loge 3/0.25
t = 4.39 hours
t = 4hrs 23mins.
Hope this helps.
:-)